40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 27 17:14:08 2015, ericq, Update, LSC, X Aux Laser crystal temperature changed 
    Reply  Fri May 29 02:05:08 2015, ericq, Update, LSC, End Laser temperatures set 
Message ID: 11328     Entry time: Wed May 27 17:14:08 2015     Reply to this: 11335
Author: ericq 
Type: Update 
Category: LSC 
Subject: X Aux Laser crystal temperature changed 

Rana suspects that the lack of X beatnote is related to the PSL laser temperature change (ELOG 11294). 

I used the information on the wiki and old elogs (wiki-40mELOG 6732), to deduce that the new end laser temperatures should be:

  • X end-> 38.98 C
  • Y end-> 35.80 C

I went out to the X end and found the laser crystal temperature set to 40.87, which is not what the measurements I linked to suggest would be the ideal temperature for the previous NPRO laser temperature of 30.89, which would be 37.02. I could not find any elog describing the choice of this setpoint.  

I've changed the X end laser crystal temperature to the value above. I've hooked up the X IR and green beatnotes to go the control room analzyer, and have been looking for the beatnote as I adjust the digital temperature offset, but haven't found it yet...

If this proves totally fruitless, we can just put the lasers back to their original temperatures, since it's unclear if it helped the PC drive noise levels. 

ELOG V3.1.3-